OpenStudy (anonymous):

how do u solve thid problem 7r^2-14r=-7

OpenStudy (aric200):

Do you need the answer or how to solve it?

OpenStudy (ranga):

7r^2 - 14r = -7 This is a quadratic equation. Make the right hand side zero by adding 7 to both sides: 7r^2 - 14r + 7 = 0 Factor 7 out 7(r^2 - 2r + 1) = 0 which implies r^2 - 2r + 1 = 0 Can you solve this quadratic equation by factoring?